if i made the flower bed 2 feet wider and 3 feet longer then it would have been 64 square feet larger. if i made it 3 feet wider and 2 feet longer it would have been 68 square feet larger. find the length and width of the flower bed.
Answer by stanbon(75887)

if i made the flower bed 2 feet wider and 3 feet longer then it would have been 64 square feet larger. if i made it 3 feet wider and 2 feet longer it would have been 68 square feet larger. find the length and width of the flower bed.
------------------
Let width = W
Let length be L
------------------------
Equations:
(W+2)(L+3)-WL = 64
(W+3)(L+2)-WL = 68
--------------------------
Modify:
3W + 2L = 58
2W + 3L = 62
-------------------
Modify:
6W + 4L = 116
6W + 9L = 186
-----------------------
Subtract and solve for "L":
5L = 70
L = 14
---
Solve for W:
2W + 3L = 62
2W + 42 = 62
2W = 20
W = 10
--------
Original dimensions: Length = 14 ft ; Width = 10 ft.
=====================================
Cheers,
Stan H.
============